# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ 4.1 Setup and Baselines
Backbone. We utilize LLaMA 3.2-1B (Dubey et al., 2024) as the backbone, a single linear layer as the projector, and MoleculeSTM (Mustafa et al., 2022) as the graph encoder for processing molecular graphs. For MoE expansion, we set $l_{\text{moe}}=1/4L$ and number of experts to 3. More details about model implementation can be found in Appendix C.
@@
Training Details. We use PyTorch (Paszke et al., 2019) with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) for more efficient parallel training. For unified tuning, we train 15 epochs with LoRA rank of 64. For separate tuning, Omni-Mol is trained for 10 epochs with the same LoRA configuration. The learning rate is set to 8e-5 from grid search for all experiments. For experiment consistency, random seed is set to 0. More details can be found in Appendix D.
@@ 4.2 Main Results
Figure 4: Scaling trend of Omni-Mol. (Left) The scaling trend respect to dataset proportion, metrics are averaged across tasks, (Right) The scaling trend respect to model size, the metrics of Quantum Mechanics Property Prediction task are normalized. we observed a clear log scaling behavior.
Model Type #Param HOMO LUMO GAP Avg.
Quantum Mechanics Property Prediction Task
Alpaca1 In-Context 6.7B - - - 322.109
LLaMA22 In-Context 6.7B 0.7367 0.8641 0.5152 0.7510
Vicuna2 In-Context 13B 0.7135 3.6807 1.5407 1.9783
Mol-Instruction Specialist 6.7B 0.0210 0.0210 0.0203 0.0210
HIGHT Specialist 6.7B 0.0056 0.0065 0.0077 0.0066
InstructMol Specialist 6.7B 0.0048 0.0050 0.0061 0.0050
Omni-Mol Generalist 1.7B 0.0047 0.0056 0.0060 0.0052
Table 3: Main results of Quantum Mechanics Property Prediction task, 1: In-Context Learning results from Fang et al. (2024), 2: 5-shot In-Context Learning results from Cao et al. (2025).
Here, we obatin the answer that Omni-Mol can achieve the best performance across almost all tasks. As the results shown in Table 1, we have the following observations. Omni-Mol significantly outperforms all specialist baselines while utilizing only 25% of the parameters. Furthermore, Omni-Mol surpasses the corresponding state-of-the-art generalist baseline by an average of approximately 1%, 4%, 13%, 15%, and 40% across forward prediction, retrosynthesis, reagent prediction, molcap, and Description Q&A separately. That is to say, Omni-Mol achieves superior performance with greater parameter efficiency, demonstrating its effectiveness in becoming a general AI chemist. Due to the page limit, we only report 6 tasks, the remaining results can be found in Appendix E.
@@ 4.3 Is Omni-Mol a Scalable Framework?
One critical property of LLMs is their scaling behavior in relation to both model and data size. In this study, we demonstrate that Omni-Mol is a scalable framework by conducting three distinct types of scaling experiments: (1) We select three different sizes of LLMs from the LLaMA 3 series, 1B, 3B, and 8B, for language backbone scaling. (2) We evaluate the impact of dataset size by down-sampling the original dataset to 20%, 40%, 60%, and 100% of its full size. (3) To examine task scaling, we train Omni-Mol on different numbers of tasks, specifically, 1, 2, 4, 8, and 15 tasks, and observe the performance of individual tasks within these multi-task settings.
(1) As shown in the left of Figure 4, we observe a clear logarithmic scaling trend as the dataset proportion increases. The relationship between the model’s average performance and the dataset proportion can be approximately expressed as $y=0.07\cdot\log(x)+0.41$, indicating that the model’s performance improves as the amount of data increases. The overlaid radar charts further demonstrate that this trend holds true across all tasks.
(2) As shown on the right side of Figure 4, the performance of Omni-Mol across all tasks increases as the model size grows. We also observe a clear logarithmic scaling trend, where this relationship can be approximated as $\text{perf}=0.02\cdot\log(x)+0.32$. However, as seen in the radar chart, the performance does not consistently improve across all tasks with the increase in model size.
(3) The average results are shown in Figure 2 based on reagent prediction. Although there is no clear elementary functional relationship between the model’s average performance and the number of tasks, Omni-Mol continues to benefit as the number of tasks increases. This indirectly supports our theory about the value of a high-quality, universal representation space. In contrast, InstructMol experiences a performance decline when the number of tasks exceeds eight, as conflicts between tasks hinder its ability to simultaneously learn all tasks effectively.
Figure 5: Ablation study of active learning-based dataset selection. Omni-Mol w. RS: we randomly sample all datasets and mix them, Omni-Mol w. AD: we use all data samples.
Figure 6: Ablation studies: (Left) Ablation of unified training. The performance is averaged across all metrics. (Middle) Ablation of adaptive gradient stabilization. (Right) Ablation of MoE expansion.
@@ 4.4 Is Unified Instruction Tuning Essential?
One key aspect of Omni-Mol is its ability to leverage unified learning across omni-molecular tasks, enabling the convergence to more generalizable representations. To evaluate this, we compare the performance of separate tuning on individual tasks (Omni-Mol w/o UT) against our unified tuning. As shown in Figure 6, Omni-Mol w/o UT performs significantly worse across five tasks compared to ours. This indicates that the representations learned through unified tuning are superior and benefit from shared knowledge. Interestingly, even tasks that are not directly related, such as the molcap task, which is distinct from both reaction and regression tasks, still show improvements with unified tuning. Additional ablation results are provided in Appendix F.
4.5 Is Dataset Selection Essential?
In Figure 5, we report the results compared with random sampling (Omni-Mol w. RS) and all-data training (Omni-Mol w. AD). The figure also includes the ratio of full data to down-sampled data. With 40% of the data, the Omni-Mol significantly outperforms (Omni-Mol w. RS) and surpasses (Omni-Mol w. AD) on the MolCap, Description QA, and Reagent tasks. This supports the hypothesis that redundant samples exist within the data, and selecting appropriate subsets from different datasets can lead to better performance. In the remaining two tasks, our results are also comparable, considering that we used only a small portion of the data.
4.6 Can We Mitigate the Conflict Collapse?
How Do Adaptive Gradient Stabilization Helps? We compare our Omni-Mol with Omni-Mol w/o ST, which replaces our adaptive module with the standard LoRA adapter. As shown in the middle of Figure 6, Omni-Mol w/o ST consistently exhibits lower performances than Omni-Mol across all tasks. This consistent decline underscores the effectiveness of our adaptive module in enhancing performance by mitigating task conflicts. This mitigation of task conflicts ensures that Omni-Mol can leverage shared knowledge without detrimental interference, thereby enhancing its ability to generalize across various tasks and modalities.
Is MoE Expansion Essential? We conduct an ablation study by replacing our Anchor-and-Reconcile Experts Expansion with a single activated Feed-Forward Network (FFN). The comparison results are shown in the right of Figure 6. We observe that Omni-Mol consistently outperforms the Omni-Mol w. FFN across all tasks, including molcap, forward prediction, retrosynthesis, and reagent prediction. Additionally, for the homo-lumo, where lower values are preferable, Omni-Mol achieves a better score. This demonstrates that Omni-Mol effectively enhances performance by leveraging specialized experts, as opposed to a single FFN. The most significant improvement is observed in forward prediction, where the diverse experts contribute to better generalization and representation learning.
@@ 4.7 Convergence Analysis via Mutual Similarity
Figure 7: Demonstration of similarity scores heatmap for methods trained on varying numbers of tasks. Surprisingly, Omni-Mol shows rising similarity scores as the task count increases (direction of the black dashed line). Compared to the decreasing trend of InstructMol, this suggests that Omni-Mol converges toward a more consistent representation space universally.
In this section, we aim to validate the theory in Section 3.6 and verify the convergence of Omni-Mol, we compute the representation sequence with model trained on 1, 2, 4, and 8 tasks, we use mutual_knn (Huh et al., 2024) as our similarity function $\Gamma(\cdot,\cdot)$, the results are shown in Figure 7.
Obviously, when the number of tasks increases, the similarity of the representations learned by Omni-Mol also increases. This indicates that the model’s representations are gradually converging. This outcome supports our earlier hypothesis that adding more tasks reduces the size of the model’s general solution space. As a result, the model is forced to learn representations within a smaller and more focused space, leading to the convergence of the learned representations. Ultimately, these representations converge to a universal form that can effectively solve all tasks.
Interestingly, in the mutual similarity analysis of InstructMol, we observe the opposite trend. As the number of tasks increases, the representations learned by InstructMol become progressively less similar to those learned previously. This suggests that with each added task, the changes in the solutions learned by InstructMol become larger, indicating that it is unable to converge to a universal representation space through unified training. In fact, the model may be moving further away from such a space.
@@ 5 Conclusion
We introduce Omni-Mol, a model that unifies 15 tasks, resolves the conflict collapse problem, and learns generalizable representations. Omni-Mol achieves this through unified SFT, MoE expansion, and active learning-based data selection. Omni-Mol achieves sota performance across multiple tasks, and we demonstrated its scalability and ability to scale up performance as the number of tasks increases. Finally, we provide experimental evidence showing that Omni-Mol achieves a more general convergent solution space, acquiring the general capability to solve diverse tasks.
Limitation. We identified two limitations: (1) Due to the limited computational resources, we are unable to further scale up the model with higher computational resources, which prevents us from exploring the limits of model’s performance. (2) Omni-Mol’s tasks are still primarily focused on molecular understanding, and it is not yet capable of designing new molecules. Future work should explore unifying understanding and generation.
Acknowledgement
The authors acknowledge the University of Maryland Institute for Advanced Computer Studies (UMIACS) for providing computational resources that contributed to this research.
Impact Statement

### v3

@@ 6.1 Experimental Setup
Backbone. We utilize LLaMA 3.2-1B dubey2024llama as the backbone, a single linear layer as the projector, and MoleculeSTM mustafa2022multimodal as the graph encoder. For MoGE expansion, we set $l_{\text{MoGE}}=1/4L$ and the number of experts to 5, there are 2 routed experts and 1 shared expert in total. More details about model implementation can be found in Appendix C.
Evaluation Metric. Following cao-etal-2024-presto , we evaluate Mol2Num tasks with MAE and $R^{2}$. For Mol2Text we adopt the standard NLP suite, which are BLEU-2, BLEU-4, ROUGE-2, ROUGE-L, and METEOR outlined in cao2023instructmol ; li2024towards . Text2Mol and reaction-related Mol2Mol tasks, following cao2023instructmol ; zhang2025atomas are gauged with Exact Match, Levenshtein score, MACCS similarity, Morgan similarity, and RDK similarity, quantifying how 1D molecular strings encode functional and structural information. For the molecule editing task in Mol2Mol, we report both unconstrained and constrained success rates. The constrained variant requires (i) the optimized molecule to satisfy the property thresholds, which are QED (Quantitative Estimate of Drug-likeness) bickerton2012quantifying $\geq 0.6$ and DRD2 score (probability of being an active DRD2 ligand) olivecrona2017molecular $\geq 0.5$, and (ii) a 2D topological similarity to its precursor of at least $\Delta=0.4$. QED is computed with RDKit. All of our evaluation metrics take into account both the linguistic quality and their biological relevance. Details are in Appendix B.
@@
Training Details. We use PyTorch paszke2019pytorch with DeepSpeed ZeRO-2 rajbhandari2020zero . For unified tuning, we train 15 epochs with GAL rank of 64. For separate tuning, model is trained for 10 epochs with the same GAL configuration. The learning rate is set to 8e-5 from the grid search for all experiments. For consistency, the random seed is set to 0. More details can be found in Appendix D.
@@ 6.2 Main Results
Figure 4: Scaling trend of Omni-Mol. (Left) The scaling trend with respect to dataset proportion, metrics are averaged across tasks. (Right) The scaling trend with respect to model size.
Model #Par Exa BLEU Lev RDK MAC Mor Val
IUPAC2SELFIES
Llama3.2-1B* SL 1.2B 0.31 0.947 16.63 0.639 0.818 0.60 0.995
DeepSeek ICL 685B 0.00 0.828 30.37 0.177 0.399 0.14 0.893
Omni-Mol GL 2.2B 0.39 0.952 13.38 0.729 0.871 0.69 0.996
Model #Param Exa BLEU Lev RDK MAC Mor Val
Text Guided Molecule Generation
Llama3.2-1B* SL 1.2B 0.04 0.789 28.11 0.447 0.629 0.329 0.941
DeepSeek ICL 685B 0.02 0.658 35.27 0.217 0.398 0.170 0.608
Omni-Mol GL 2.2B 0.12 0.824 23.59 0.562 0.721 0.442 0.963
Table 4: Results of Text2Mol tasks, * means we train the model with LoRA and the same multimodal configuration as Omni-Mol.
Figure 5: Ablation studies: (Left) Averaged performance comparison, the results demonstrate that InstructMol with joint training outperforms separate training. Further integrating MoGE yields additional performance gains across all tasks. (Middle) Ablation of Gradient Adaptive LoRA. (Right) Ablation of MoGE expansion. For the mid and right figure, the left axis represents the average performance of Reagent, Molcap and MolDesign. The right axis represents the average performance of Yield Prediction.
Here, we obtain the answer that Omni-Mol can achieve the best performance across almost all tasks. As the results shown in Table 3, we have the following observations. Omni-Mol significantly outperforms almost all specialist baselines while utilizing only 33% of the parameters. Furthermore, Omni-Mol surpasses the corresponding state-of-the-art generalist baseline by an average of approximately 5%, 7%, 9%, 11%, and 40% across forward prediction, retrosynthesis, reagent prediction, molcap, and Description Q&A separately. We further notice that on Mol2Num tasks, Omni-Mol improves the $R^{2}$ score by $21.4\%$, and lowers the Mean Absolute Error(MAE) by $25.1\%$, $25.8\%$, $39.3\%$ on Weight, LogP and TPSA regression respectively. That is to say, Omni-Mol achieves superior performance with greater parameter efficiency, demonstrating its effectiveness in becoming a general AI chemist. To further examine GAL’s adaptability to 3D molecular tasks, we report Omni-Mol’s results on ten tasks. Omni-Mol achieves the best performance on all ten tasks, outperforming the above baselines. Additional experimental details are provided in Appendix E.
@@ 6.3 Is Omni-Mol a Scalable Framework?
One critical property of LLMs is their scaling behavior in relation to both model and data size. In this study, we demonstrate that Omni-Mol is a scalable framework by conducting two distinct types of scaling experiments: (1) We select three different sizes of LLaMA 3 series, 1B, 3B, and 8B, for language backbone scaling. More backbone results are in Appendix F. (2) We evaluate the impact of dataset size by down-sampling the original dataset to 20%, 40%, 60%, and 100% of its full size.
(1) Data Scaling. As shown in the left of Figure 4, we observe a clear scaling trend as the dataset proportion increases, indicating that the model’s performance improves as the amount of data increases. This suggests that further increasing the dataset size can bring more benefits and build a stronger and more generalized chemical AI.
(2) Parameter Scaling. As shown on the right side of Figure 4, the performance of Omni-Mol across all tasks increases as the model size grows. We also observe a clear scaling trend. However, this trend is not as pronounced as the performance gains resulting from increased dataset size, which suggests that there remains potential for further expansion in the amount of data.
@@ 6.4 Is Unified Instruction Tuning Essential?
We evaluate whether our Omni-Mol dataset can enhance unified instruction tuning. We select one task from each of the following categories: Mol2Mol, Mol2Text, Mol2Num, and Text2Mol, and retrain InstructMol on each. We compare the results of individually trained models (one LoRA per task) with those trained in a unified manner (a single LoRA shared across all tasks), as shown in Figure 5. The results indicate that unified learning on the Omni-Mol dataset yields performance improvements. Interestingly, the molcap and yield tasks exhibit noticeable differences from the other tasks, yet still benefit from unified fine-tuning, with yield regression showing the greatest performance gain.
6.5 Ablations on MoGE
How do MoGE help? Building upon the unified training of InstructMol on our dataset, we incorporate MoGE into the framework. As illustrated in Figure 5, InstructMol with Joint Training and MoGE consistently outperforms both InstructMol with Separate Training and InstructMol with Joint Training, demonstrating the effectiveness of MoGE. By adapting to the intrinsic dimensionality of different tasks and leveraging the specialization among MoGE experts, MoGE enhances Omni-Mol’s ability to generalize across a diverse range of tasks and modalities.
Is GAL essential? We compare our Omni-Mol with Omni-Mol w/o GAL, which replaces the Gradient Adaptive LoRA with the standard LoRA. As shown in the middle of Figure 5, Omni-Mol w/o GAL consistently exhibits lower performances than Omni-Mol. This consistent decline underscores the effectiveness of our GAL in enhancing performance by adaptively adjusting itself to the intrinsic dimension.
Is MoGE expansion essential? We conduct an ablation study by removing the MoGE expansion and training the model with GAL alone (Exclude MoE). The comparison results are shown on the right side of Figure 5. We observe that Omni-Mol consistently outperforms the Omni-Mol w/o MoE across multiple tasks, including reagent prediction, molcap, yield regression, and text-guided molecule generation. These results demonstrate that Omni-Mol effectively enhances performance by leveraging specialized experts. The most significant improvement is observed in yield regression, where the diversity of experts contributes to better generalization and representation learning.
@@ 6.6 Convergence Analysis via Mutual Similarity
Omni-Mol is trained across a wide range of tasks, we aim to examine how its learned representations vary with respect to the number of tasks involved in training.
If the model is indeed learning within a progressively smaller solution space, we should observe a convergence in representation similarity. This is because, under a larger number of tasks, the reduction in the solution space is expected to constrain the variation in the learned representations. We compute the representation sequence with the model trained on 1, 2, 4, and 8 tasks with
Figure 6: Demonstration of similarity scores heatmap for methods trained on varying numbers of tasks.
mutual_knn huh2024platonic , the results are shown in Figure 6. Obviously, when the number of tasks increases, the similarity of the representations learned by Omni-Mol also increases. This indicates that the model’s representations are gradually converging. This outcome supports the hypothesis that adding more tasks reduces the size of the model’s general solution space. Ultimately, these representations converge to a universal form that can effectively solve all tasks. Interestingly, in the mutual similarity analysis of InstructMol, we observe the opposite trend. As the number of tasks increases, the representations learned by InstructMol become progressively less similar to those learned previously. This suggests that with each added task, the changes in the solutions learned by InstructMol become larger, indicating that it is unable to converge to a universal representation space through unified training.
@@ 7 Conclusion
We introduce Omni-Mol, a model that unifies 16 tasks and Omni-Mol dataset, with over 1.4 million samples. Omni-Mol learns generalizable representations and achieves this through unified tuning, MoGE expansion, and Gradient Adaptive LoRA. Omni-Mol achieves SOTA performance across multiple tasks, and we demonstrate its scalability and ability to scale up performance as the number of tasks increases. Finally, we provide experimental evidence showing that Omni-Mol achieves a more general convergent solution space, acquiring the general capability to solve diverse tasks.
Limitations
We identified two limitations: (1) Due to the limited computational resources, we are unable to further scale up the model with higher computational resources, which prevents us from exploring the limits of model’s performance. (2) Omni-Mol’s tasks are still primarily focused on small molecules, future work should explore proteins and the interaction between proteins and small molecules.
Broader Impacts